L. Collect data from several fast food chains on the number of fat calories and grams of saturated fat in menu items. Record at least 12 ordered pairs of (fat calories, grams of saturated fat) Organize your data in a table. Il. Make a scatter plot of the data on graph paper. Be sure to label the axes and use an appropriate title for the graph. You may wish to use a graphing calculator, spread sheet, or other technology resource (such as the graphing utility link below) to aid you in graphing Create a Graph II Draw a trend line for the scatter plot. Use the following scatter plot of the ordered pairs (fat grams, total calories) as an example 600 550 500 450 400 350 300 250 200 150 100 50 1020 304 Fat Grams IV. Calculate the slope of the trend line. (Choose two points on the line and find vertical change over horizontal change.)
V. Using the slope and y-intercept, write the equation of the trend line (y mx+b). VI. Choose a "calories from fat" value that is not in your collected data set and that is at least 10 fat calories away from any collected value. Use the equation calculated in step V to predict the number of fat grams in an item having that number of fat calories. Be sure to show your work VII. Search for an item in a fast food menu having the same number of fat calories as the one you chose above. (If you cannot find the exact value, get as close as you can.) Compare the calculated value from step VI to this actual value. Explain why (or why not) you would have expected your prediction (calculated value) to be close to the actual value.
Answer #1

if we extend the regression line and the x-axis as given in the following image the

(v)regression line =y=mx+b=10x+125

since the regression line touches x-axis at -10 so b=10 and it touches y-axis at 125 so m=125

(vi) for x=1, y=10*1+125=135

(vii) for x=12, y=10*12+125=245,

actual value of y=225

10 20 30 40 Fat Grams